OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, –3) and has a slope of 1/2?

OpenStudy (anonymous):

@Eyad could u help me again

OpenStudy (anonymous):

2x + y = 4 2x – y = 4 x + 2y = 4 x – 2y = 4

OpenStudy (anonymous):

sure, Use the law : y2-y1=m(x2-x1) where (-2,-3)=(x1,y1) ,and m=slope.. can u do it now ?

OpenStudy (anonymous):

this is where im stuck

OpenStudy (anonymous):

The point–slope form of the equation for a line passing through the point (x1,y1) is given by: Assuming (-2, –3) = (x1, y1) and m = 172 and substituting the values, we get: Y-Y1=m(X-X1) Y-(-3)=1/2(x-(-2)) Y+3=1/2(X+2) Y+3=1/2X+1 Y=1/2X-2

OpenStudy (anonymous):

m=1/2 not 172 lol , the answer is correct , as pointed above

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

OK Then: y-(-3)=1/2(x-(-2)) y+3=1/2x+1 y=1/2x-2----------> (multiply by 2) 2y=x-4 Or in another form x – 2y = 4 ____ Got it ?

OpenStudy (anonymous):

not really but thanks anyway

OpenStudy (anonymous):

what part you didn't understand ,I just fill up the rule ?

OpenStudy (anonymous):

what seems to be the problem , everything is clearly written ---

OpenStudy (anonymous):

how did u get 172?

OpenStudy (anonymous):

its not 172 ,, i meant to write 1/2

OpenStudy (anonymous):

ohh ok i get it then

OpenStudy (anonymous):

ok .